OpenStudy (anonymous):

To prove: 3(sin^4(3pi/2-Alpha)+Sin^4(3pi+Alpha)) -2(sin^6(pi/2+Alpha) sin^6(5Pi-Alpha)) =2

OpenStudy (anonymous):

maybe use equation editor so its easier to read?

OpenStudy (anonymous):

\[3\left[ \sin ^{4}\left( 3\pi \div2+\alpha \right) +\sin ^{4}\left( 3\pi+\alpha \right)\right]-2\left[ \sin^6\left( \pi/2+\alpha \right) \sin^6\left( 5\pi-\alpha \right)\right]=1\]